OpenStudy (anonymous):

What is the solution set of |x – 6| = 4?

OpenStudy (diyadiya):

|x-6| = 4 there are two solutions: x-6 = 4 x-6=-4 add 6 on both sides Add 6 on both sides x = 10 x = 2

OpenStudy (anonymous):

What is the solution set of |x + 3| – 2 = –5?

OpenStudy (diyadiya):

can you try that on your own ?

OpenStudy (anonymous):

i did and i got –5 or 5 ANDi got it wrong can you show me how to do it

OpenStudy (diyadiya):

|x + 3| – 2 = –5 Add 2 on both sides |x+3| = -3 But absolute value can't be negative so there are no solutions

OpenStudy (anonymous):

What is the solution set of |2x + 4| < 12? is x > 8 and x < 4 correct

OpenStudy (diyadiya):

Yes! that's right :)

OpenStudy (anonymous):

yesssssssssssss XD

OpenStudy (anonymous):

What is the solution set of |x| > 12? i got x < –12 or x > 12

OpenStudy (diyadiya):

Yes!! :)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form i am confused